OpenStudy (anonymous):

divide and simplify. (21z^3v - 5z^5v^3 + 24z^7v^5) / (-3z^5v^3)

OpenStudy (anonymous):

i have to simplify it further and i'm having a hard time

OpenStudy (anonymous):

@iGreen. @YOLO_KING

OpenStudy (anonymous):

Divide all the whole numbers by 3. 24 / 3 = ? -5/3 21 / 3 = ?

OpenStudy (anonymous):

8, 1.7, 7

OpenStudy (anonymous):

Yes, now we subtract all the exponents of 'v' by 2 4 - 2 = ? 2 - 2 = ?

OpenStudy (anonymous):

mmk so 2 and 0

OpenStudy (anonymous):

ugh somehow I still got it wrong this is my last section of questions and I can never seem to fully get it right ughhh the answer came out looking really odd D=

OpenStudy (anonymous):

Hmm.. It should've been: \(8v^2z^2 - 1.7 + 7\) Add: \(8v^2z^2 + 5.3\)